# p-MoD: Building Mixture-of-Depths MLLMs via Progressive Ratio Decay

### v1

@@ C.1 Detailed Training Configurations
In this section, we provide our detailed training configurations, as shown in Table 7. LLaVA-1.5 employs a fixed input resolution of 336$\times$336, while LLaVA-NeXT supports a pre-defined set of different resolutions (up to 672$\times$672). Both the LLaVA-v1.5 models and the LLaVA-NeXT models go through the same pre-training stage, where the MLP connector module is trained on 558K image caption data [25] with a learning rate of 1e-3 and a batch size of 256.
@@
During supervised fine-tuning, LLaVA-1.5 is trained on 665K instruction-tuning data [23], while LLaVA-NeXT is trained on 779K data [^1] https://huggingface.co/datasets/lmms-lab/LLaVA-NeXT-Data. The vision encoder of the LLaVA-NeXT baseline model is updated for fine-tuning. For our p-MoD-LLaVA-NeXT model, we freeze the vision encoder to save training time, as the available GPU resources are limited. When measuring the training GPU hours reported in Table 3, we freeze the vision encoder for both baseline and p-MoD models to ensure fair comparison.

### v2

@@ D.1 Detailed Training Recipe
Our detailed training recipe of p-MoD models is shown in Table 12. LLaVA-1.5 employs a fixed input resolution of $\times336$, while LLaVA-NeXT supports a pre-defined set of different resolutions (up to $\times672$). Both the LLaVA-v1.5 models and the LLaVA-NeXT models go through the same pre-training stage, where the MLP connector module is trained on 558K image caption data [29] with a learning rate of 1e-3 and a batch size of 256.
@@
During supervised fine-tuning, LLaVA-1.5 is trained on 665K instruction-tuning data [27], while LLaVA-NeXT is trained on a larger set of 779K data***https://huggingface.co/datasets/lmms-lab/LLaVA-NeXT-Data. The vision encoder of the LLaVA-NeXT baseline model is updated for fine-tuning. For our p-MoD-LLaVA-NeXT model, we freeze the vision encoder to save training time, as the available GPU resources are limited. When measuring the training GPU hours reported in Table 7, we freeze the vision encoder for both baseline and p-MoD models to ensure fair comparison.
D.2 Hardware and Hyperparameters
We train all our models on 8 NVIDIA RTX A6000 GPUs. The inference efficiency metrics reported in Section 4.5 are measured on a single A6000 GPU. By default, we set the gating factor $\alpha$ in TanhNorm to 0.2, and the shift factor $\beta$ in PRD to $0.5$.
Due to computational constraints, we mainly use 7B models in our experiments. Experiments on 13B models are show in Table 8 and Section A.1.
D.3 Details on the PRD schedule.
We find it challenging for MoD layers to learn to predict meaningful weights when the token retention ratio is set to an extremely large or small value. To address this issue, we constrain the range of the token retention ratio $R_{l}$ within a predefined maximum and minimum threshold. If $R_{l}$ exceeds the maximum threshold, we set the token retention ratio to $1$ so that the MoD module is not applied to the $l$-th layer. If $R_{l}$ falls below the minimum threshold, the token retention ratio for the $l$-th layer is set to the minimum threshold:
$R_{l}^{\prime}=\begin{cases}1,&\text{if}\quad R_{l}\geq max\\ R_{l},&\text{if}\quad min<R_{l}<max\\ min,&\text{if}\quad R_{l}\leq min\end{cases}\quad.$ (6)